# Ordinary primes in Hilbert modular varieties

### v1

@@ 1 Introduction
Let $\mathcal{X}$ be a projective smooth scheme over the ring $\mathcal{O}_{F,S}$ of $S$-integers in a number field $F$, and let $X/F$ be the generic fibre. For each integer $i$, one defines the Hodge polygon $HP^{i}(X)$ from the Hodge filtration on $H^{i}_{dR}(X/F)$, and, for every $\mathfrak{p}$ outside the finite set $S$, the Newton polygon $\mathrm{NP}(\operatorname{Frob}_{\mathfrak{p}},H^{i}_{\mathrm{cris}}((\mathcal{X}\otimes_{\mathcal{O}_{F}}k(\mathfrak{p}))/W(k(\mathfrak{p}))))$ using the crystalline Frobenius.
Katz conjectured, and Mazur proved [20], that the latter lies above the former:
@@
When the two coincide, we say that $\mathfrak{p}$ is an ordinary prime (in degree $i$ for $X$), following Mazur. The following appears to have been first considered by Serre [28, no. 133] for abelian varieties $X$ (in degree $1$):
Conjecture 1.1 (Ordinariness Conjecture)
@@
It is known to be true for (elliptic curves and) abelian surfaces by an argument of Katz, and for abelian varieties whose endomorphism ring is $\mathbb{Z}$ and whose algebraic monodromy group satisfies a condition: See Pink [23, §7]. It is also known for all CM abelian varieties.
@@
More generally, we consider the parts of the intersection cohomology of the Hilbert modular varieties attached to totally real number fields $F$ of degree $d=[F:\mathbb{Q}]$, cut out by new normalised cusp forms $f$ of parallel weight $(2,\cdots,2)$. We note that ordinariness in this context appeared as an assumption in the construction of Galois representations by Wiles [33], which was later removed by Taylor [29] and by Blasius and Rogawski [5].
@@
We show that these conjectures (Conjectures 2.3.1 and 2.3.2) satisfy a basic consistency, by proving that the Hodge-Tate polygon does not depend on the choice of a prime $\lambda$ of $F$ lying over $\ell$, and that the conjectures are independent of the prime $\ell$, by using theorems of Gabber, of Katz and Laumon, of André and of de Cataldo and Migliorini.
Let us return to the Hilbert modular varieties and the forms $f$. The Fourier coefficients of $f$ generate a number field $K_{f}$, which is either totally real or CM, through which the Hecke correspendences act on the part $M(f)$ cut out by $f$ and all its conjugates. The conjectures make sense and are independent of $\ell$ for $M(f)$.
@@
By using the construction of the Galois representation attached to $f$ and the purity of $IH$, we first show that $M(f)$ satisfies the analogue of the Katz Conjecture and that we can push the Newton polygon ‘half way’ to ordinariness in a quantifiable sense, for a principally abundant set of primes. However, in the attempt to push just beyond the half-way threshold, we face an obstruction of “geometry of numbers” type (Minkowski). We show how to overcome it (for a principally abundant set of primes) by using a stronger form of Sato-Tate equidistribution (see §3.3), but this last remains unknown in general.
In order to go further and to obtain unconditional results, we look into (a) ‘multivariate’ variants of the Sato-Tate Conjecture in §3.3 and (b) the interaction between $F$ and $K_{f}$ in §3.5. For the latter, we define an invariant, the slope $\sigma_{F}(K)\in[0,1]$ of a coefficient number field $K$ over a ground number field $F$ (see Definition 3.5.4), by using the action of $\operatorname{Gal}(\overline{\mathbb{Q}}/F)$ on the set $\operatorname{Hom}(K,\overline{\mathbb{Q}})$ of field embeddings. The slope $\sigma_{F}(K)$ is 00 if, for example, (1) $[K:\mathbb{Q}]$ is a prime not dividing $[F:\mathbb{Q}]$; or (2) the Galois group of $K/\mathbb{Q}$ is the (full) symmetric group on $[K:\mathbb{Q}]$ letters111this ‘Maeda-like’ condition appears to be often satisfied in practice, but not always. See Section 5 for examples. and $K$ and $F$ have coprime discriminants.
@@
    $[K_{f}^{\circ}:\mathbb{Q}]\leq 2$; where $K_{f}^{\circ}$ is the smallest Frobenius field of $f$ (à la Ribet);
@@
In Section 5, we look at the forms $f$ with low levels for $4$ number fields $F$ of degree $d\leq 4$, and show that for most $f$ under consideration, (a), (b) and (d) provide an abundant set of ordinary primes unconditionally, and that (c) complements them under the Sato-Tate condition. We give descriptions of the conditional cases as well as some cases where our methods fall short of yielding abundance of ordinary primes.
Nonconstant coefficients
In the final Section 6, we formulate analogues of the Katz and the ordinariness conjectures for submotives of the intersection cohomology of more general motivic coefficients, following a suggestion of Katz.
In cases where we have good crystalline realisations compatible with the $\ell$-adic realisations (which include the nonconstant motivic coefficients on Hilbert modular varieties), we verify the Katz conjecture by using Mazur’s theorem. In case the submotive has potentially abelian $\ell$-adic realisation, we also verify the ordinariness conjecture by using Serre’s theory [26]. Finally, we provide some methods to deal with the parallel motivic weight $(3,\cdots,3)$ in the Hilbert modular case.
@@ 2.1 Polygons
Definition 2.1.1
Assume that $\rho$ is $\mathbb{Q}$-rational in the sense of Serre [26].
@@
Equivalently: Choose an isomorphism ¯¯$\overline{\mathbb{Q}}_{\ell}\simeq\overline{\mathbb{Q}}_{p}$ and let $x_{1},\cdots,x_{m}\in\overline{\mathbb{Q}}_{\ell}$ be the eigenvalues of $\operatorname{Frob}_{\mathfrak{p}}$ on $V$, and the multiset of slopes
@@
There seem to be competing sign conventions for the Hodge-Tate weights. We take the ‘geometric’ one, so that $H^{1}$ of an elliptic curve has the Hodge-Tate weights $\{0,1\}$.
@@ 2.2 Independence of Hodge-Tate weights
First, we show that the Hodge-Tate polygon we have defined coincides with the classical Hodge polygon for all André motives, thereby extending a theorem of Faltings [12], [13].
Theorem 2.2.1
@@
Then the set with multiplicities of the Hodge-Tate weights of $M_{p}$ (suitably normalised) coincides with that of the complex Hodge numbers of $M_{\sigma}$.
@@
This follows from the definition of André motivated cycles, together with the fact that the comparison isomorphisms in display are isomorphisms between Weil cohomology theories (and as such compatible with pullback, pushforward, cup product, cycle class, and Poincaré duality, that are involved in the definition of André motivated cycles). See André [1, §2.3 and §2.4].
@@
Let $\pi:Y\longrightarrow X$ be any resolution of singularities. We use the main theorem of de Cataldo and Migliorini [9], fortified in $K$-rationality by Patrikis [21, §8] to deduce the existence of an André motivated cycle
@@
acting on $H^{n}_{dR}(Y)$. Clearly, there is a comparison isomorphism of $IH^{n}_{dR}(X/K)\otimes_{K}\mathbb{C}$ with the (transcendental) Hodge structure on the Betti realisation $IH^{n}(\sigma(X),\mathbb{Q})\otimes\mathbb{C}$. Moreover, this last Hodge structure coincides with the Hodge structure contsructed by Morihiko Saito, see de Cataldo [8, Th. 4.3.5].
@@
Strictly speaking, the results on the intersection cohomology can be proven by using the de Cataldo-Migliorini theorem [9] only (and not using the $K$-rational version [21]). To see this, note that the Hodge numbers and the Hodge-Tate weights are insensitive to the base change to a finite extension of $K$ (in both local and global cases), and the construction of [9] yields the necessary André motivated cycle over a finite extension of $K$.
Remark 2.2.6
Here is a variant for the Baily-Borel compactification $X^{BB}$ of a Shimura variety $X$ defined over the reflex field $E$:
One has the action of the Hecke algebra (with $\mathbb{Q}$-coefficients) on the André intersection motive $\mathfrak{ih}^{n}(X^{BB})$, and one can cut out the Hecke-isotypic components $M$, as André motives.
A sketch of an argument that does not leave the category of (pure) André motives: The finite correspondences on $X$ extend canonically to $X^{BB}$. While the same is not true of any single fixed toroidal compactification in general, the extension holds true in the ‘tower’ of toroidal compactifications in which ever finer cone decompositions are allowed, see Pink [22, Prop. 6.25].
One then checks that the action of such correspondences preserves the direct summands in the cohomology of toroidal compactifications coming from the intersection cohomology of $X^{BB}$. This way one gets the action of the Hecke algebra (with $\mathbb{Q}$-coefficients) on the André intersection motive via André motivated cycles.
This fact, that the $\mathbb{Q}$-Hecke isotypic components of $\mathfrak{ih}^{n}(X^{BB})$ are André motives, also follows from an unpublished result of F. Ivorra and S. Morel, which deals with more general Nori motives.
@@
All this applies in particular to the Hilbert modular varieties and the submotive $M(f)$ of $\mathfrak{ih}^{d}(X^{BB})$ cut out by any new cuspform $f$ of parallel weight $(2,\cdots,2)$ and its conjugates, and Corollary 2.2.3 applies to it.

### v3

# To Professor Richard Taylor with gratitude

@@ Abstract
11F41 (primary), 14G35, 11F30, 11G18 (secondary).
keywords:
Hilbert modular form, ordinary reduction, Sato-Tate equidistribution.
@@ 1 Introduction
Let $\mathcal{X}$ be a projective smooth scheme over the ring $\mathcal{O}_{F,S}$ of $S$-integers in a number field $F$, and let $X/F$ be the generic fibre. For each integer $i$, one defines the Hodge polygon $HP^{i}(X)$ from the Hodge filtration on $H^{i}_{dR}(X/F)$: it is the convex planar graph emanating from the origin in which the line segment of slope $a$ appears with the multiplicity equal to the Hodge number $h^{a,i-a}=\dim_{F}H^{i-a}(X,\Omega^{a})$. For every $\mathfrak{p}$ outside the finite set $S$, one also forms the Newton polygon $\mathrm{NP}(\operatorname{Frob}_{\mathfrak{p}},H^{i}_{\mathrm{cris}}((\mathcal{X}\otimes_{\mathcal{O}_{F}}k(\mathfrak{p}))/W(k(\mathfrak{p}))))$ using the $p$-adic slopes of the crystalline Frobenius. (See §3.4.)
Katz [K71, Conjecture 2.9] conjectured, and Mazur proved [Ma73], that the latter lies above the former:
@@
When the two coincide, we say that $\mathfrak{p}$ is an ordinary prime (in degree $i$ for $X$), following Mazur. The following appears to have been first considered by Serre [Se13, no. 133] for abelian varieties $X$ (in degree $1$):
Conjecture 1.1 ((Ordinariness Conjecture)).
@@
[…]
@@
More generally, we consider the parts of the intersection cohomology of the Hilbert modular varieties attached to totally real number fields $F$ of degree $d=[F:\mathbb{Q}]$, cut out by new normalised cusp forms $f$ of parallel weight $(2,\cdots,2)$. The ordinariness in this context first appeared as an assumption in the construction of Galois representations by Wiles [Wi88], which was later removed by Taylor [T89] and by Blasius and Rogawski [BR93]. However, many of the best results and constructions in Iwasawa and Hida theory at present depend on the existence of ordinary primes in a crucial way: See among others the works of Emerton, Pollack and Weston [EPW06]; of Ochiai [Oc06]; of Nekovář [N01]; of Dimitrov [Di13]; of Skinner and Urban [SkUr14]; and of Wan [Wn15].
@@
These conjectures (Conjectures 2.4 and 2.5) satisfy a basic consistency, in that (a) the Hodge-Tate polygon is independent of the choice of a prime $\lambda$ of $F$ lying over $\ell$; and (b) the conjectures are independent of the auxiliary prime $\ell$. We prove these statements by using theorems of Gabber (on the independence of $\ell$ in the intersection cohomology of complete varieties) [Fu00], of Katz and Laumon (on the constructibility properties of certain constructions in derived categories) [KL85], of André (his theory of motivated cycles) [A96] and of de Cataldo and Migliorini (on the motivated nature of the decomposition theorem in intersection cohomology) [dCM15].
Let us return to the Hilbert modular varieties and the forms $f$. By using either (i) the theorem of de Cataldo and Migliorini in op. cit. and its rational extension due to Patrikis [Pa16] or (ii) the recent motivic constructions of Ivorra and Morel [IM19], we construct an intersection motive of $X$ which in realisations give the intersection cohomology. Then by lifting the action of the Hecke correspondences on the intersection cohomology to one on the intersection motive (see Proposition 2.2.6), we construct an André motive $M(f)$. The conjectures make sense for these submotives, and the consistency mentioned above also hold for them. We denote by $K_{f}$ the Hecke field of $f$.
@@
By using the construction of the Galois representation attached to $f$ and the purity of $IH$, we first show that $M(f)$ satisfies the analogue of the Katz Conjecture and that we can push the Newton polygon ‘half way’ to ordinariness in a quantifiable sense, for a principally abundant set of primes. However, in the attempt to push just beyond the half-way threshold, we face an obstruction of “geometry of numbers” type (Minkowski). We show how to overcome it (for a principally abundant set of primes) by using a stronger form of Sato-Tate equidistribution (see §3.3), but this last form remains unknown in general.
In order to go further and to obtain unconditional results, we look into (a) ‘multivariate’ variants of the Sato-Tate Conjecture in §3.3 and (b) the interaction between $F$ and $K_{f}$ in §3.5. For the latter, we define an invariant, the slope $\sigma_{F}(K)\in[0,1]$ of a coefficient number field $K$ over a ground number field $F$ (see Definition 3.5.4), by using the action of $\operatorname{Gal}(\overline{\mathbb{Q}}/F)$ on the set $\operatorname{Hom}(K,\overline{\mathbb{Q}})$ of field embeddings. The slope $\sigma_{F}(K)$ is 00 if, for example, (1) $[K:\mathbb{Q}]$ is a prime not dividing $[F:\mathbb{Q}]$; or (2) the Galois group of $K/\mathbb{Q}$ is the (full) symmetric group on $[K:\mathbb{Q}]$ letters111this ‘Maeda-like’ condition appears to be often satisfied for the Hecke fields $K=K_{f}$ in practice, but not always. See Section 5 for examples. and $K$ and $F$ have coprime discriminants.
@@
    $[K_{f}^{\circ}:\mathbb{Q}]\leq 2$; where $K_{f}^{\circ}$ is the smallest Frobenius field of $f$ in the sense of Ribet (see §3.2 for the precise definition);
@@
In Section 5, we look at the forms $f$ with low levels for $4$ number fields $F$ of degree $d\leq 4$, and show that for most $f$ under consideration, (a), (b) and (c) provide an abundant set of ordinary primes unconditionally, and that (d) complements them under the strong Sato-Tate condition. We give descriptions of the conditional cases as well as some cases where our methods fall short of yielding abundance of ordinary primes.
Plan. Here are some main ingredients and ideas in the text.
In Section 2, we show (Theorem 2.2.1) that for André motives, the Hodge polygon made of the Hodge numbers in its (transcendental) Betti realisations coincide with the Hodge-Tate polygon made of the $p$-adic Hodge-Tate weights in the (algebraic) $p$-adic étale realisations. This applies in particular to the Hecke isotypic components of the intersection cohomology (motive) of the Hilbert modular varieties.
In Section 3, we introduce a few things in preparation for the Main Theorem in Section 4. (a) In §3.4 we study the partially ordered semiring of Newton polygons; this will be useful in dealing with the tensor induction in Theorem 4.1.1. (b) In §3.5, we define the notion of slope and bisection in the interaction between the ground field $F$ and the coefficient field $K$; very roughly speaking, they measure the sizes and shapes of ‘large’ orbits in the Hecke and Frobenius fields. In Theorem 4.1.1, these will be combined with the Chebotarev density theorem and strong forms of the Sato-Tate Conjecture.
In Section 4 we prove the Main Theorem 4.1.1. Here we make a connection (which to the author’s knowledge is new) between two conjectures on the eigenvalues of the Frobenius elements: Namely, the Sato-Tate Conjecture on the archimedean properties of the eigenvalues (which in turn is closely linked to the Langlands Program) on the one hand and the Ordinariness Conjecture on the $p$-adic properties (with varying $p$) of the eigenvalues on the other hand.
In the final Section 6, we formulate analogues of the Katz and the ordinariness conjectures for submotives of the intersection cohomology of more general motivic coefficients, following a suggestion of Katz. In cases where we have good crystalline realisations compatible with the $\ell$-adic realisations (which include the nonconstant motivic coefficients on Hilbert modular varieties), we verify the Katz conjecture by using Mazur’s theorem. In case the submotive has potentially abelian $\ell$-adic realisation, we also verify the ordinariness conjecture by using Serre’s theory [Se98]. Finally, we provide some methods to deal with the parallel motivic weight $(3,\cdots,3)$ in the Hilbert modular case.
@@ 2.1 Polygons
Definition 2.1.1.
Assume that $\rho$ is $\mathbb{Q}$-rational in the sense of Serre [Se98].
@@
Equivalently: Choose an isomorphism ¯¯$\overline{\mathbb{Q}}_{\ell}\simeq\overline{\mathbb{Q}}_{p}$ and let $x_{1},\cdots,x_{m}\in\overline{\mathbb{Q}}_{\ell}$ be the eigenvalues of $\operatorname{Frob}_{\mathfrak{p}}$ on $V$. Then, the multiset of slopes
@@
There seem to be competing sign conventions for the Hodge-Tate weights. We take the ‘geometric’ one, so that $H^{1}$ of an elliptic curve has Hodge-Tate weights $\{0,1\}$.
@@ 2.2 Independence of Hodge-Tate weights
First, we show that the Hodge-Tate polygon we have defined coincides with the classical Hodge polygon for all André motives, thereby extending a theorem of Faltings [Fa88], [Fa89]. For this we will crucially rely on the theory of motivated cycles and the resulting category of André motives, given in [A96].
Theorem 2.2.1.
@@
Then the set with multiplicities of the Hodge-Tate weights of $M_{p}$ coincides with that of the complex Hodge numbers of $M_{\sigma}$.
@@
This follows from the definition of André motivated cycles, together with the fact that the comparison isomorphisms in display are isomorphisms between Weil cohomology theories (and as such compatible with pullback, pushforward, cup product, cycle class, and Poincaré duality, that are involved in the definition of André motivated cycles). See André [A96, §2.3 and §2.4].
@@
Let $\pi:Y\longrightarrow X$ be any resolution of singularities. We use the main theorem of de Cataldo and Migliorini [dCM15], strengthened in $K$-rationality by Patrikis [Pa16, §8] to deduce the existence of an André motivated cycle
@@
acting on $H^{n}_{dR}(Y)$. Clearly, there is a comparison isomorphism of $IH^{n}_{dR}(X/K)\otimes_{K}\mathbb{C}$ with the (transcendental) Hodge structure on the Betti realisation $IH^{n}(\sigma(X),\mathbb{Q})\otimes\mathbb{C}$. Moreover, this last Hodge structure coincides with the Hodge structure contsructed by Morihiko Saito, see de Cataldo [dC12, Th. 4.3.5].
@@
Strictly speaking, the results on the intersection cohomology can be proven by using the de Cataldo-Migliorini theorem [dCM15] only (and not using the $K$-rational version [Pa16]). To see this, note that the Hodge numbers and the Hodge-Tate weights are insensitive to the base change to a finite extension of $K$ (in both local and global cases), and the construction of [dCM15] yields the necessary André motivated cycle over a finite extension of $K$.
In case $X=X^{BB}$ is the Baily-Borel compactification of a Shimura variety $X^{\circ}$ (we refer to Ash-Mumford-Rapoport-Tang [AMRT] and Pink [Pi98] in general, and Brylinski-Labesse [BL84] and Rapoport [Ra78] in the special case of Hilbert modular varieties), one further decomposes the intersection cohomology of $X$ into the Hecke-isotypic components: the Hecke correspondences act on the intersection cohomology of $X$, and span a $\mathbb{Q}$-subalgebra $\mathcal{H}_{X,\mathbb{Q}}$ in the finite dimensional $\mathbb{Q}$-algebra $\operatorname{End}_{HS}(IH^{d}_{B}(X,\mathbb{Q}))$. By decomposing $\mathcal{H}_{X,\mathbb{Q}}$ into a product of $\mathbb{Q}$-algebras, we obtain the Hecke isotypic components.
Proposition 2.2.6.
The Hecke isotypic components come from André motives over the reflex field $E$.
As a consequence, Theorem 2.2.1 applies to these components.
Proof.
This boils down to first finding an André (pure Nori) motive $\mathfrak{ih}(X)$ whose $\ell$-adic and Betti realisations give the $\ell$-adic and Betti intersection cohomology of $X$; and then lifting the action of the Hecke correspondences to one on $\mathfrak{ih}(X)$.
While the first step can be done as above in an ad hoc fashion — using a (noncanonical) resolution of singularities and using the theorems of de Cataldo, Migliorini and Patrikis — the second step is done most systematically (in our opinion) by using the theory of weight filtration. We learned the argument from S. Morel (cf. [Mo08, §5]), which uses the more recent motivic constructions of Ivorra and Morel [IM19].
(Before the details, let us stress the main point and indicate where the innovations are. The intersection complex $IC$ (as a perverse sheaf in the derived category of constructible sheaves) was originally constructed as an iterated application of the $2$-step procedure: taking the direct image $Rj_{\ast}$ under open immersions $j$, and then truncating with respect to a topological stratification and a function called ‘perversity’. See the explicit formula [BBD, Prop. 2.1.11].
One of the innovations in [Mo08] was to realise $IC$ (for the middle perversity) over finite fields as the truncation with respect to the weight filtration: See [Mo08, Th. 3.1.4].
This renders the extension of the Hecke operators to $IC$ deceptively easy: One needs neither to worry about singularities in the boundary (which can be bad) — to which one must pay heed if one uses the original (topological) definition — nor to rely on the toroidal compactifications — of which there is no canonical choice, and many are needed to extend the Hecke correspondences. This is why we adopt the idea.
In [IM19], Ivorra and Morel construct the four operations of Grothendieck (namely $f^{\ast}_{\mathcal{M}}$, $f^{!}_{\mathcal{M}}$, $f_{\ast}^{\mathcal{M}}$, $f_{!}^{\mathcal{M}}$ for morphisms $f$ between quasiprojective varieties) and the weight filtrations on the derived category of ‘perverse mixed motives’; this last abelian category is moreover shown to be equivalent to the Nori motives.
This allows one to formally apply the algorithm described in [Mo08, §5], but this time applied in the motivic derived category of [IM19], rather than in the derived category of constructible sheaves as in [Mo08]. Then the Betti and $\ell$-adic realisations of these constructions give rise to the Hecke operators constructed previously on the level of complexes of sheaves.)
Now let us turn to a more detailed argument, and indicate which constructions in [IM19] replace those in the parallel argument from [Mo08, §5].
Let $j:X^{\circ}\hookrightarrow X=X^{BB}$ be the open immersion of the Shimura variety and $c_{1},c_{2}:Y^{\circ}\longrightarrow X^{\circ}$ be the two finite étale maps that give rise to a given Hecke correspondence. With $j^{\prime}:Y^{\circ}\hookrightarrow Y$ denoting the open immersion into the Baily-Borel compactification, the $c_{i}$ extend canonically to $\overline{c_{i}}:Y\longrightarrow X$. We start with the identity correspondence
$u=1:c_{1}^{\ast}\mathbb{Q}_{X^{\circ}}\longrightarrow c_{2}^{!}\mathbb{Q}_{X^{\circ}}.$
arising from the natural isomorphisms $c_{1}^{\ast}\mathbb{Q}_{X^{\circ}}\simeq\mathbb{Q}_{Y^{\circ}}$ and $c_{2}^{!}\mathbb{Q}_{X^{\circ}}\simeq\mathbb{Q}_{Y^{\circ}}$.
Here and below, the functors $f^{\ast}$ and $f^{!}$, etc., refer to the ones in [IM19, Th. 5.1], where the notations $f^{\ast}_{\mathcal{M}}$, $f^{!}_{\mathcal{M}}$, etc., are used.
Just as in [Mo08, §5.1], but using the motivic constructions of the $4$ functors and the base change morphisms, stated [IM19, Th. 5.1], we take $Rj^{\prime}_{\ast}={j^{\prime}}_{\ast}^{\mathcal{M}}$ and use the base change morphisms
${\overline{c_{1}}^{\ast}Rj_{\ast}\mathbb{Q}_{X^{\circ}}\longrightarrow Rj^{\prime}_{\ast}c_{1}^{\ast}\mathbb{Q}_{X^{\circ}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{u}$${Rj^{\prime}_{\ast}c_{2}^{!}\mathbb{Q}_{X^{\circ}}\longrightarrow\overline{c_{2}}^{!}Rj_{\ast}\mathbb{Q}_{X^{\circ}}.}$
We then use the fact that the lowest weight filtration of $Rj_{\ast}\mathbb{Q}_{X^{\circ}}$ is canonically isomorphic to the intersection complex $j_{!\ast}(\mathbb{Q}_{X^{\circ}}[d])[-d]$. For this, it is enough to show that the (motivic) weight filtration has $\ell$-adic realisation equal to the ($\ell$-adic) weight filtration. But this follows from the definitions and construction: [IM19, Def. 6.12, Def. 6.13, and Prop. 6.16].
Then we proceed as in [Mo08, Lem. 5.1.4] to obtain
$\overline{u}:\overline{c_{1}}^{\ast}IC_{X}\longrightarrow\overline{c_{2}}^{!}IC_{X},$
lifting the cohomological correspondence in realisations. Again the point is that all the arrows in display in [Mo08, Lem. 5.1.4] are constructed using (only) the functoriality of $Rj_{\ast}$, the base change morphisms, and the weight filtration. In our (motivic) context, we use the main theorem [IM19, Th. 5.1] for the first two and the weight filtration constructed in [IM19, Prop. 6.16] for the last. ∎
@@
This applies in particular to the Hilbert modular varieties and the submotive $M(f)$ of $\mathfrak{ih}^{d}(X^{BB})$ cut out by any new cuspform $f$ of parallel weight $(2,\cdots,2)$ and its conjugates, and Corollary 2.2.3 applies to it.